OpenStudy (anonymous):

Help on an essay question. Part 1: Solve each of the quadratic equations below. Show your work. x2 − 16 = 0 and x2 = −2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

I already solved the first equation I just can't solve the second one.

OpenStudy (anonymous):

Well for the second equation make it equal 0 and make sure your a is positive. \[x^{2} = −2x + 24\] Subtract \[x^{2}\] On both sides. \[0 = -x^{2}-2x+24\] Multiply both sides by -1 \[0=x^{2}+2x-24\] Now can you solve it?

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^{2}-4ac}}{ 2a }\]

OpenStudy (anonymous):

Still can't solve it? I feel so stupid. I got as far as you got with the multiplying it by -1 but can't go any further?

OpenStudy (anonymous):

Well your a=1 b=2 c=-24

OpenStudy (anonymous):

But what do I do with that?

OpenStudy (anonymous):

\[\frac{ -2 \pm \sqrt{2^{2}-4(1)(-24)} }{ 2(1) }\] \[\frac{ 2 \pm \sqrt{4-(-96)} }{ 2 }\] \[\frac{ 2 \pm \sqrt{100} }{ }\] \[\frac{ 2 \pm 10 }{ 2 }\]

OpenStudy (anonymous):

So \[\frac{ 2+10 }{ 2 }=6\] \[\frac{ 2-10 }{ 2 }=-4\] So x = 6, and x = -4

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

No. sorry. I am only in the ninth grade doing algebra.

OpenStudy (anonymous):

You're supposed to take Algebra I in 9th grade.

OpenStudy (anonymous):

So for the equation you have \[x^{2}+2x-24=0\] So this is what it would look like if I wrote the parent function. \[ax^{2}+bx+c=0\] To solve the quadratic equation you use this, the quadratic formula. \[x=\frac{ -b \pm \sqrt{b^{2}-4ac}}{ 2a }\] So for our equation \[x^{2}+2x-24=0\] a=1 b=2 c=-24 So when we plug it in to the quadratic formula (as I showed you above it shows you what x equals.

OpenStudy (anonymous):

Making a little more sense to you?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

So x = 1?

OpenStudy (anonymous):

No I solved it for you x=6 and x=-4

OpenStudy (anonymous):

Well I have to go to a class now. So if you need more help ask jim_thompson5910

OpenStudy (anonymous):

But a is a negative 1

OpenStudy (anonymous):

a can't be negative. As I have said above. Well for the second equation make it equal 0 and make sure your a is positive. x^2=−2x+24 Subtract x^2 on both sides. 0=−x2−2x+24 Multiply both sides by -1 0=x^2+2x−24 Now put the 0 on the other end. x^2+2x−24=0 Now your a is positive and your b is now positive and your c is negative.

OpenStudy (anonymous):

I got it now thank you